# Calibration of a $\Delta$E-E telescope based on CeBr3 scintillator for secondary charged particles measurements in hadron therapy

### v1

@@ Abstract
Hadrontherapy is a promising cancer treatment method that offers better dose conformity and reduces damage to healthy tissues compared to conventional radiotherapy. However, one major challenge remaining is the precise characterization of secondary particles generated by nuclear interactions of the primary beam with tissues. Current data on secondary charged particles, such as protons and light ions, remain insufficient, particularly in the clinically relevant energy ranges. This lack of experimental data introduces uncertainties in treatment planning softwares and Monte Carlo calculations, thus compromising the accuracy of dose delivery to the patients.
This work consists in the characterization of secondary charged particles generated in hadrontherapy using a $\Delta$E-E telescope comprising a CeBr3 crystal scintillator and a plastic scintillator. The calibration and response of this telescope to ions commonly used in clinical settings is presented in this work, highlighting adherence to Birks’ law for accurate energy measurements.
This study is the first to optimize a $\Delta$E-E telescope combining CeBr3 and plastic scintillators specifically for secondary particle detection in hadrontherapy. This represents an important step in the exploitation of the system for nuclear data acquisition, as it enables both the measurement of energy and the discrimination of secondary particles. The objective is to develop a system compatible with clinical use, allowing for the most precise possible comparison with treatment planning software calculations.
@@ 1 Introduction
However, the interactions between the primary particle beam and human tissue lead to nuclear fragmentation, producing secondary particles such as lighter ions, neutrons, and gamma rays. These secondary particles can cause various types of damage, including ionization, excitation, and fragmentation of molecular bonds. This damage can affect cellular functions and lead to biological responses that can contribute to undesired dose in healthy organs and tissues [29] [9].
Accurate dose calculations for particle therapy are critically dependent on high-performance algorithms that model complex physical and biological processes, often using data generated through Monte Carlo simulations [28] [3]. However, there is currently a lack of experimental data regarding nuclear reactions in particle therapy, which can introduce inaccuracies in dose calculations [19][25].
@@
One key point of the CLINM project is to measure the yields, charges, and energies of the fragments produced under conditions that replicate clinical experimental settings. In this work, the focus will be on the charged secondary particles. To achieve this, a $\Delta$E-E telescope detection system has been developed. This technique leverages the principles of energy loss and remaining energy to discriminate different particle charges. The thinner detector in the telescope, a plastic scintillator, measures the energy loss ($\Delta$E) of charged particles as they pass through, while a CeBr3 crystal scintillator measures the remaining energy (E) of the particle.
@@
In this study, the primary beams simulate the clinical energy ranges used in therapy, but the focus remains on the secondary charged particles generated during beam interactions with materials mimicking patient tissues (e.g., PMMA). This telescope is designed to measure energy deposition patterns of these secondary charged particles, which include protons, light ions (e.g., deuterons, alpha particles), and heavier fragments generated by nuclear fragmentation processes, with energies ranging from a few MeV up to 200 MeV/u.
@@ 2.1 Detectors description
The $\Delta$E-E telescope is composed of two scintillating detectors: a thin plastic scintillator ($\Delta$E) and a thick crystal scintillator of CeBr3 (E), long enough to stop 12C up to 200 MeV/u (Figure 2).
The CeBr3 detector is a cylindrical 2$\times$2 inches crystal scintillator, from Advatech UK Ltd [1], coupled to a R6231-100 Hamamatsu photomultiplier (PMT). An entrance window in front of the crystal inside the CeBr3 is made of 400 $\mu$m of aluminum and 1 mm of polytetrafluoroethylene (PTFE), used as a reflector. The CeBr3 detector has a short decay time of 19 ns, a high light yield of 60,000 Photons/MeV, an energy resolution of 3.8 keV at 662 keV fo X-ray and gamma-ray, and an inherent low background radiation of 0.004 Bq/cm3.
@@
Two different photomultipliers (PMT) from Hamamatsu, R7057 and XP3990, were tested, coupled to the plastic scintillator, supplied with -1100V and -1200V high-voltage respectively, in order to optimize the settings, i.e. minimize saturation while maximizing resolution.
Both detectors signals were aquired in coincidence by a WaveCatcher digitizer module, developed by the LAL laboratory (Paris, France) [6]. The sampling frequency was set to be 3.2 GHz on the DAQ. Figure 1 illustrates digitized signals obtained from the plastic scintillator and CeBr3 crystal when irradiated with 200 MeV/u 12C ions. As expected, the pulse duration and the rising time of the plastic scintillator ($\sim$ 1.4 ns) are shorter than the ones from the CeBr3 ($\sim$ 19 ns), in accordance with intrinsic properties of both scintillators. A gate of 80 ns was applied for the detector’s coincidence detection.
@@ 2.2 Experimental setup
Experiments were carried out in different facilities providing ions of various types and energies: Cyrcé cyclotron at IPHC (Strasbourg, France) [12], producing protons beam of 25MeV, which can be attenuated by an aluminium wheel to lower energies ; CAL (Centre Antoine Lacassagne, Nice, France) protontherapy center [17], specializes in proton therapy treatments for cancer, providing a proton beam of 62 MeV ; cave M of GSI (Darmstadt, Germany), the heavy ion accelerator facility SIS-18 [14], using PMMA thickness variations to modulate carbon beam energies ; CNAO hadrontherapy center (Pavia, Italy) [23], which offers advanced hadron therapy treatments using carbon ions and protons ; the LISE spectrometer [4] of the GANIL facility (Caen, France), producing lithium ion beam of 63.6 MeV/u and carbon ion beam of 75.4 MeV/u, which can be degraded by PMMA wheel to lower energies. The characteristics of the beams which were used are summarised in Table 1. The different experimental setups for each facilities are presented on Figure 2.
Facility Ion type Energy range
Cyrcé 1H 16 - 25 MeV
CAL 1H 61 MeV
GSI 12C 120 - 180 MeV/u
CNAO 1H 80 - 180 MeV
CNAO 12C 120 - 200 MeV/u
GANIL 12C 30 - 75 MeV/u
GANIL 8Li 45 - 64 MeV/u
@@
At Cyrcé, aluminum attenuators from 0.147 mm to 1.424 mm were used to degrade the beam energy from 25 MeV down to 16 MeV, while at GSI, three different PMMA thicknesses were used in order to achieve carbon beam energy ranges of 120 MeV/u, 160 MeV/u and 180 MeV/u. At CAL and CNAO, the energy of the beam was not attenuated. Only one energy of 61 MeV/u was used at CAL, while several energies of 12C and protons were used at CNAO, from 120 MeV/u to 200 MeV/u. In the experimental setup at GSI, a plastic scintillator of 1 mm was placed in front of the beam for another experiment which was carried out at the same beam time. At GANIL, to attenuate the energy of the beam, a specific experimental setup with a wheel was developed and build, with thirty different PMMA thicknesses in order to achieve carbon beams from 30 MeV/u to 64 MeV/u and lithium beams from 45 MeV/u to 75 MeV/u. The plastic detector was replaced with two silicon detectors, each 0.2 mm thick, for an additional test phase aimed at improving resolution for the CLINM project.
@@ 2.3 Analysis of the data
The used analysis software package was an extension of the QAPIVI software tool [11], called STIVI (Software for Tracking of Ions and Vertex Imaging). The baseline refers to the initial portion of the waveform where no significant signal is expected, representing the underlying noise level of the system. To establish the baseline of the waveforms, the projection of the first 10 bins in time was employed on the pulse shape. Subsequently, the mean value and dispersion, indicative of the noise level, were computed for accurate characterization.
The determination of the signal’s rise and fall involved setting thresholds relative to the baseline, typically defined as the baseline minus 10 times the noise level.
The integration of the signal giving the charge was then performed within the determined range, while the amplitude of the signal was identified as its maximum value. The baseline of the waveforms was subtracted to the signal, to isolate the signal from the background noise. The distributions of the amplitudes of the Waveforms (in mV) were fitted with Gaussian functions, in order to extract the mean amplitude value and its corresponding standard deviation.
@@
$\frac{dL}{dx}=\frac{S\cdot\frac{dE}{dx}}{1+k_{B}\cdot\frac{dE}{dx}}$
@@
$A=\frac{S^{\prime}\cdot E+A_{0}}{1+k_{B}^{\prime}\cdot E}$
where $A$ is the amplitude of the light signal measured (in mV), $E$ is the total energy deposited (in MeV), $S^{\prime}$ is a conversion factor between the deposited energy and the amplitude, $k_{B}^{\prime}$ is a modified quenching parameter adapted to total energy (in MeV-1), $A_{0}$ represents a pedestal or baseline signal (in mV) accounting for background noise or residual detector response.
The adjusted conversion factor $S^{\prime}$ serves as a proportionality factor to convert the deposited energy $E$ into the expected amplitude $A$, reflecting both light yield efficiency and the detector’s response, with typical values ranging from $\sim 10^{-2}$ to $10^{0}$ mV.MeV-1 [18][5][8]. Unlike $S$, which relates to differential energy loss, $S^{\prime}$ is calibrated to reflect the detector’s pulse response to total energy deposition. The adjusted quenching parameter $k_{B}^{\prime}$ represents saturation effects related to total energy deposited, with values typically ranging from $10^{-2}$ to $10^{-3}\,\text{MeV}^{-1}$ for inorganic scintillators [10] and from $10^{-3}$ to $10^{-4}\ \text{MeV}^{-1}$ for organic or plastic scintillators [7], aligning with Birk’s quenching model. The pedestal $A_{0}$ accounts for baseline signal amplitude from residual detector response, typically in the low mV range. The adapted Birk’s law maintains the quenching effect via $k_{B}^{\prime}$ while incorporating linear responses at low energy deposition through $S^{\prime}$.
For the next sections, the function will be used as follows, with $p_{0}$, $p_{1}$ and $p_{2}$ the variable parameters, and, for simplicity, $S^{\prime}$ and $k_{B}^{\prime}$ will be refereed to as $S$ and $k_{B}$, respectively:
$A=\frac{p_{0}*E+p_{1}}{1+p_{2}*E}$ (2.1)
These three parameters are dependent on the geometry of the scintillators and the high voltage applied to the PMT. Thus, they were considered as free parameters in the fit of the detector response to ions, and represent the calibration parameters values of our detectors.
@@ 2.4 Monte Carlo simulations
The Monte Carlo code Geant4 10.07 [3] with the INCL++ physic list [2] was used to evaluate the deposited energy in the plastic scintillator, and the energy of the ions reaching the CeBr3 crystal, after its entrance window. While it is not necessary to take into account the thickness of this entrance window for $\gamma$-rays detection, it needs to be considered for ions detection, as they will lose a non-negligible amount of energy inside.
@@ 3.1 Response of the plastic scintillator
The energy resolution of the scintillator with the XP3990 PMT is also shown, as a function of the deposited energy, on Figure 4(b). This analysis integrates considerations of experimental beam straggling effects, which may introduce uncertainties in the measured energy resolution. Indeed, the energy deposited by the ions can exhibit significant variation due to straggling, leading to a spread in the measured energy spectrum. The uncertainty in energy loss can be in the order of several MeV, depending on the thickness and composition of the material. Similarly, scattering can cause angular deviations up to a few degrees, impacting the spatial resolution of the detector.
@@ 3.2 Response of the CeBr3 to ions
Figure 5: Calibration curves of the CeBr3 crystal scintillator obtained with both protons and 12C, with an applied voltage of + 350 V for (a) and obtained with protons, 8Li and 12C, with an applied voltage of + 400 V for (b). Energies can be found on Table 1.
All the fit parameters values collected for both plastic and CeBr3 and for all the voltage used can be found in Table 2. It can be observed that, depending on the voltage and the photomultiplier gain, distinct Birk’s constants are obtained. Nevertheless, it is noteworthy that these discrepancies in Birk’s constants across different conditions overlap within the margin of error and fall within the expected range reported in the literature, thereby indicating coherence in the data.
Detector S kB (MeV-1)
Plastic -1200V 8.4 $\pm$ 0.8 0.0074 $\pm$ 0.0026
Plastic -1100V 2.30 $\pm$ 0.16 0.0048 $\pm$ 0.0019

### v2

@@ Abstract
Hadrontherapy is an established cancer treatment method that enables a more localized dose deposition compared to conventional radiotherapy, potentially reducing the dose to surrounding healthy tissues in certain clinical cases. However, a key limitation in current treatment planning lies in the limited experimental data available for the characterization of secondary particles generated by nuclear interactions of the primary beam with tissues, which directly impacts the accuracy of Monte Carlo tools and analytical models used in dose calculations. Indeed, this leads to the adoption of larger safety margins and can limit the use of hadrontherapy for treating certain complex or sensitive tumor locations.
This work is part of the context of the characterization of secondary charged particles generated by ion beams in the energy range relevant for particle therapy applications, using a $\Delta E-E$ telescope comprising a CeBr3 crystal scintillator and a plastic scintillator. The calibration and response of this telescope to ions commonly used in clinical settings is presented in this work, highlighting adherence to Birks’ law for accurate energy measurements.
This study is the first to optimize a $\Delta E-E$ telescope combining CeBr3 and plastic scintillators specifically for secondary particle detection in hadrontherapy. It represents an essential step toward the experimental acquisition of nuclear data, enabling accurate measurement and identification of secondary charged particles generated by therapeutic beams in tissue-equivalent materials. The system is designed for use in controlled experimental setups that reproduce clinical conditions, with the goal of improving the predictive accuracy of treatment planning software through enhanced Monte Carlo simulation inputs.
@@ 1 Introduction
However, the interactions between the primary particle beam and human tissue lead to beam and target fragmentation, producing secondary particles such as lighter ions, neutrons, and gamma rays. These secondary particles can contribute to out-of-field dose deposition through ionization and nuclear interactions in surrounding tissues. This damage can affect cellular functions and lead to biological responses that can contribute to undesired dose in healthy organs and tissues [3] [4]. This additional dose is not accounted for by the primary beam alone and must be carefully modeled to ensure accurate treatment planning and risk assessment.
Accurate dose calculations for particle therapy rely primarily on analytical or semi-empirical models integrated in treatment planning systems, that model complex physical and biological processes. These models are often supported or benchmarked using data from detailed Monte Carlo simulations such as Geant4 [5] [6]. However, there is currently a lack of experimental data regarding nuclear reactions in particle therapy, which can introduce inaccuracies in dose calculations [7][8], potentially affecting clinical treatment margins and robustness, as highlighted in [9].
@@
One key point of the CLINM project is to measure the yields, charges, and energies of the fragments produced under conditions that replicate clinical experimental settings, using tissue-equivalent materials. In this work, the focus will be on the charged secondary particles. To achieve this, a $\Delta E-E$ telescope detection system has been developed. This technique leverages the principles of energy loss and remaining energy to discriminate different particle charges. The thinner detector in the telescope, a plastic scintillator, measures the energy loss ($\Delta E$) of charged particles as they pass through, while a CeBr3 crystal scintillator measures the remaining energy ($E$) of the particle.
@@
In this study, the primary beams simulate the clinical energy ranges used in therapy, but the focus remains on the secondary charged particles generated during beam interactions with materials mimicking patient tissues (e.g., PMMA). This telescope is designed to measure energy deposition patterns of these secondary charged particles, which include protons, light ions (e.g., deuterons, alpha particles), and heavier fragments generated by nuclear fragmentation processes, in materials mimicking human tissues (e.g., PMMA phantoms), under experimental conditions that simulate those of clinical beams with energies ranging from a few MeV up to 200 MeV/u.
@@ 2.1 Detectors description
The $\Delta E-E$ telescope is composed of two scintillating detectors: a thin plastic scintillator ($\Delta E$) and a thick crystal scintillator of CeBr3 ($E$), long enough to stop 12C up to 200 MeV/u.
The CeBr3 detector is a cylindrical 2$\times$2 inches crystal scintillator, from Advatech UK Ltd coupled to a R6231-100 Hamamatsu photomultiplier (PMT) [10]. An entrance window in front of the crystal inside the CeBr3 is made of 400 $\mu$m of aluminum and 1 mm of polytetrafluoroethylene (PTFE), used as a reflector. The CeBr3 detector has a short decay time of 19 ns, a high light yield of 60,000 Photons/MeV, an energy resolution of 3.8 keV at 662 keV fo X-ray and gamma-ray, and an inherent low background radiation of 0.004 Bq/cm3.
@@
Two different photomultipliers (PMT) from Hamamatsu [13], R7057 and XP3990, were tested, coupled to the plastic scintillator, supplied with - 1100 V and - 1200 V high-voltage respectively, in order to optimize the settings, i.e. minimize saturation while maximizing resolution.
Both detectors signals were aquired in coincidence by a WaveCatcher digitizer module, developed by the LAL laboratory (Paris, France) [14]. The sampling frequency was set to be 3.2 GHz on the the data acquisition system (DAQ). Figure 1 illustrates digitized signals obtained from the plastic scintillator and CeBr3 crystal when irradiated with 200 MeV/u 12C ions. As expected, the pulse duration and the rising time of the plastic scintillator are shorter than the ones from the CeBr3, in accordance with intrinsic properties of both scintillators. A gate of 80 ns was applied for the detector’s coincidence detection.
@@ 2.2 Experimental setup
Experiments were carried out in different facilities providing ions of various types and energies:
  • Cyrcé cyclotron at IPHC (Strasbourg, France) [15], producing protons beam of 25 MeV, which can be degraded by an aluminium wheel to lower energies;
  • CAL (Centre Antoine Lacassagne, Nice, France) protontherapy center [16], specializes in proton therapy treatments for cancer, providing a proton beam of 62 MeV ;
  • cave M of GSI (Darmstadt, Germany), the heavy ion accelerator facility SIS-18 [17], using PMMA thickness variations to modulate carbon beam energies ;
  • CNAO hadrontherapy center (Pavia, Italy) [18], which offers advanced hadron therapy treatments using carbon ions and protons ;
  • the LISE spectrometer [19] of the GANIL facility (Caen, France), producing lithium ion beam of 63.6 MeV/u and carbon ion beam of 75.4 MeV/u, which can be degraded by PMMA wheel to lower energies.
The characteristics of the beams which were used are summarised in Table 1. The different experimental setups for each facilities are presented on Figure 2.
Facility Ion type Degrador Energy range
Cyrcé 1H aluminum from 0.147 mm to 1.424 mm 16 - 25 MeV
CAL 1H none 61 MeV
GSI 12C PMMA 120 - 180 MeV/u
CNAO 1H none 80 - 180 MeV
CNAO 12C none 120 - 200 MeV/u
GANIL 12C PMMA wheel 30 - 75 MeV/u
GANIL 8Li PMMA wheel 45 - 64 MeV/u
@@
In the experimental setup at GSI, a plastic scintillator of 1 mm was placed in front of the beam for another experiment which was carried out at the same beam time. At GANIL, to degrade the energy of the beam, a specific experimental setup with a wheel was developed and build, with thirty different PMMA thicknesses in order to achieve carbon beams from 30 MeV/u to 64 MeV/u and lithium beams from 45 MeV/u to 75 MeV/u. The plastic detector was replaced with two silicon detectors, each 0.2 mm thick, for an additional test phase aimed at improving resolution for the CLINM project.
@@ 2.3 Analysis of the data
The used analysis software package was an extension of the QAPIVI software tool [20], called STIVI (Software for Tracking of Ions and Vertex Imaging). The baseline refers to the initial portion of the waveform where no significant signal is expected, representing the underlying noise level of the system. The baseline was calculated by averaging the first 10 time bins of the signal, corresponding to a pre-signal region where no event is expected. The standard deviation of this interval was taken as the noise level. The determination of the signal’s rise and fall involved setting thresholds relative to the baseline, typically defined as the baseline minus 10 times the noise level.
The integration range was defined as the interval between the rise and fall times of the signal, while the amplitude of the signal was identified as its maximum value. For each measurement condition, the distribution of signal amplitudes (in mV) were recorded over multiple events. These distributions were then fitted with Gaussian functions to extract the average amplitude — used for the detector calibration — and the standard deviation, which provides information on the signal resolution.
@@
$\frac{dL}{dx}=\frac{S\cdot\frac{dE}{dx}}{1+k_{B}\cdot\frac{dE}{dx}}$ (2.1)
@@
$A=\frac{S^{\prime}\cdot E+A_{0}}{1+k_{B}^{\prime}\cdot E}$ (2.2)
where $A$ is the amplitude of the light signal measured (in mV), $E$ is the total energy deposited (in MeV), $S^{\prime}$ is a conversion factor between the deposited energy and the amplitude, $k_{B}^{\prime}$ is a modified quenching parameter adapted to total energy (in MeV-1), $A_{0}$ represents a pedestal or baseline signal (in mV) accounting for residual signal contributions, such as baseline shifts, electronic noise, or constant components of the detector response. While the physical expectation for ideal systems would be $A_{0}=0$, practical implementations often lead to small but non-zero pedestal values, which are accounted for in the fitting procedure.
The adjusted conversion factor $S^{\prime}$ serves as a proportionality factor to convert the deposited energy $E$ into the expected amplitude $A$, accounting for the intrinsic scintillation light yield as well as detector-specific effects such as photon detection efficiency, photomultiplier gain, and signal shaping, with typical values ranging from $\sim 10^{-2}$ to $10^{0}$ mV.MeV-1 [22][21][23]. Unlike $S$, which relates to differential energy loss, $S^{\prime}$ is calibrated to reflect the detector’s pulse response to total energy deposition. The adjusted quenching parameter $k_{B}^{\prime}$ represents saturation effects related to total energy deposited, with values typically ranging from $10^{-2}$ to $10^{-3}\,\text{MeV}^{-1}$ for inorganic scintillators [24] and from $10^{-3}$ to $10^{-4}\ \text{MeV}^{-1}$ for organic or plastic scintillators [25], aligning with Birk’s quenching model. The pedestal $A_{0}$ accounts for baseline signal amplitude from residual detector response, typically in the low mV range. The adapted Birk’s law maintains the quenching effect via $k_{B}^{\prime}$ while incorporating linear responses at low energy deposition through $S^{\prime}$. It is important to emphasize that this formulation of Birks’ law is used here as a phenomenological model to effectively parametrize the nonlinear response of the detection system. The parameters $S^{\prime}$, $k_{B}^{\prime}$, and $A_{0}$ are not intrinsic properties of the scintillating material alone, but rather effective parameters that depend on the full detection chain, including photomultiplier gain, applied voltage, and signal processing conditions. Therefore, the values obtained in this work should not be interpreted as fundamental characteristics of CeBr3, but rather as calibration parameters valid under specific experimental conditions. These three parameters, including the pedestal $A_{0}$, are dependent on the geometry of the scintillators and the high voltage applied to the PMT. Thus, they were considered as free parameters in the fit of the detector response to ions, and represent the calibration parameters values of our detectors.
@@ 2.4 Monte Carlo simulations
The Monte Carlo code Geant4 10.07 [6] with the INCL++ physic list [26] was used to evaluate the deposited energy in the plastic scintillator, and the energy of the ions reaching the CeBr3 crystal, after its entrance window. The INCL++ physics list was chosen for its proven accuracy in simulating nuclear fragmentation of light ions in the 100–400 MeV/u range, which is highly relevant to hadrontherapy scenarios.
@@ 3.1 Response of the plastic scintillator
The energy resolution of the scintillator with the XP3990 PMT is also shown, as a function of the deposited energy, on Figure 4(b). This analysis takes into account the contribution of beam straggling and scattering, as evaluated by Geant4 simulations. However, some experimental uncertainties such as beam energy spread, angular divergence, or alignment tolerances are not fully modeled and may introduce additional broadening in the measured spectra. These effects can contribute to a slight overestimation of the detector’s intrinsic resolution. The uncertainty in energy loss can be in the order of several MeV, depending on the thickness and composition of the material. Similarly, scattering can cause angular deviations up to a few degrees, impacting the spatial resolution of the detector.
@@ 3.2 Response of the CeBr3 to ions
Figure 5: Calibration curves of the CeBr3 crystal scintillator obtained with both protons and 12C, with an applied voltage of + 350 V for (a) and obtained with protons, 8Li and 12C, with an applied voltage of + 400 V for (b). Energies can be found on Table 1. The solid lines correspond to fits using Birks’ law, and the shaded bands represent the propagated uncertainties from the fit parameters.
All the fit parameters values collected for both plastic and CeBr3 and for all the voltage used can be found in Table 2. As mentioned in section 2, the values of $k_{B}^{\prime}$ and $S^{\prime}$ obtained here are effective and reflect not only the quenching behavior of the scintillator but also the influence of experimental parameters such as the applied high voltage and PMT gain. The variation in the $k_{B}^{\prime}$ values observed across different configurations (e.g., +350 V vs. +400 V) is therefore expected and does not indicate a contradiction with the underlying physics. It can be observed that, depending on the voltage and the photomultiplier gain, distinct Birk’s constants are obtained.
Although the absolute values of $k_{B}^{\prime}$ vary between ions and voltage settings, they remain within the expected order of magnitude (10-3–10-2 MeV-1) reported in the literature for inorganic scintillators. As these constants are effective and depend on experimental parameters, perfect agreement between all conditions is not expected. Instead, the observed consistency in trends and ranges supports the validity of the calibration model. For all detector configurations, the pedestal was found to be close to zero, confirming that the baseline subtraction was effective.
Detector $S^{\prime}$ $k_{B}^{\prime}$ (MeV-1)
Plastic -1100V 8.4 $\pm$ 0.8 0.0074 $\pm$ 0.0026
Plastic -1200V 2.30 $\pm$ 0.16 0.0048 $\pm$ 0.0019